4. (4 pts) Consider the surface z = z2y + y). (a) Find the normal direction of the tangent plane to the surface through (1,1, 4. (4 pts) Consider the surface z=x2y+y3.(a) Find the normal direction of the tangent plane to the surface through (1,1,2).(b) Find the equation of the tangent plane in (a).(c) Determine the value a so that the vector−→v=−−→i+ 2−→j+a−→k is parallel to the tangent plane in (a).(d) Find the equation of the tangent line to the level curve of the surface through (1,1).
Answer #1

the given Z=_x2y + ys = this define finy, z) = x²y +43 - 2 f(x, y, z)=0 gin abore sarfon of is normal to surface the o f = ofPage No. Dato C © Note n is plallal to v = -1 +2jt akan an = ( 21, , -d) = (1, 2,-2) = -1 2 this a = -x = 2 the a = 2 с © Not